OpenStudy (anonymous):

Sixty is 12% of what number?

OpenStudy (anonymous):

its 100/12 *60=500

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

breja, I have an additional question for you. Willing to solve it?

OpenStudy (anonymous):

yea sure

OpenStudy (anonymous):

50 is 30% of what number? Using the method matricked taught you or with your full understanding of this concept, find the answer.

OpenStudy (anonymous):

@OrthodoxMan To find the 50% of a number to equal 30, you divide: 30 / 0.5 = 60 50% of 60 = 30.

OpenStudy (anonymous):

Sorry but i said 30% of fifty and not 50% of 30, try again.

OpenStudy (anonymous):

oops sorry

OpenStudy (anonymous):

That's alright, take another shot.

OpenStudy (anonymous):

30 / 50% = 30*100/50 = 60

OpenStudy (anonymous):

Nope. You made the same mistake again. Don't worry. You need to find what number is multiplied by 30% to get fifty. Just as i asked, fifty is 30% of what number?

OpenStudy (anonymous):

@OrthodoxMan is it correct?

OpenStudy (anonymous):

Try again?

OpenStudy (anonymous):

I'll try

OpenStudy (anonymous):

You can do it, just remember. 30% of a number = 50. You're doing the opposite. But you got the right method.

OpenStudy (anonymous):

15?

OpenStudy (anonymous):

The number has to be greater than 50. You want me to show you how to do it?

OpenStudy (anonymous):

yes please because It's very hard

OpenStudy (anonymous):

It's not, lemme use the method matricked showed you. (100%/30%)*50 = approx 166.67

OpenStudy (anonymous):

Hence, this is the equation: (100%/percentage given)*number given = answer

OpenStudy (anonymous):

ohh okay i get it know

OpenStudy (anonymous):

Wanna try one more?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

forty is 25% of what number?

OpenStudy (anonymous):

do i multiply 40 and 25 ?

OpenStudy (anonymous):

Use the equation I have given you: (100%/percentage given)*number given = answer

OpenStudy (anonymous):

ook

OpenStudy (anonymous):

What is the percentage given?

OpenStudy (anonymous):

100% /45/*12?

OpenStudy (anonymous):

Wrong numbers. forty is 25% of what number?

OpenStudy (anonymous):

is/of = %/100 is =12 of =? or x % = 60% then 12/x =60/100 cross multiply 12 * 100=1200 60 * x = 60x 1200=60x solve for x by dividing both sides by 60 1200/60= 20!

OpenStudy (anonymous):

Is 20 your final answer?^

OpenStudy (anonymous):

@OrthodoxMan yeah.

OpenStudy (anonymous):

ITs alright everyone! I have solved everything for her, return to your remedial duties good citizens...

OpenStudy (anonymous):

25*45/100 = 10

OpenStudy (anonymous):

lmfao @EnteringStacis

OpenStudy (anonymous):

what exactly do you want from "25*45/100 = 10"?

OpenStudy (anonymous):

please expalain it a dat better

OpenStudy (anonymous):

oops 25 times 40 divide by 100 then 10

OpenStudy (anonymous):

what do you mean "then 10"?

OpenStudy (anonymous):

=10?

OpenStudy (anonymous):

(25*40)/100 = 45/4 fyi

OpenStudy (anonymous):

that doesnt make sense... are you looking if the statement is true or false?

OpenStudy (anonymous):

just tell me exactly what they ask you in the problem

OpenStudy (anonymous):

thts it.... im closing thi s question thank you and what pecent of 40 is 25

OpenStudy (anonymous):

breja, i believe someone said earlier, but all you have to do is ask yourself that question and substitute "what number" with X. So you'll come up with: X * .4 = 25 Simply solve for X and you'll find your answer.